# Quantum router of silicon-vacancy centers via a diamond waveguide

### v1

@@ Abstract
As a key component of quantum networks, quantum router distributes quantum information among different quantum nodes. Silicon-vacancy (SiV) center in diamond offers a promising platform for quantum technology due to its strong strain-induced coupling with phonons. However, the development of practical quantum router faces challenges of achieving long-range entanglement and suppressing decoherence. Here, we propose a non-Markovian quantum router based on a diamond waveguide embedded with an array of SiV centers as the quantum nodes. Unlike conventional channel-switching methods, our design enables parallel quantum-state transfer from a single input node to multiple target nodes, analogous to a classical WiFi router. We demonstrate that persistent entanglement and suppressed decoherence of the SiV centers over long distances are achievable when bound states are present in the energy spectrum of the total system formed by the SiV centers and the phonon waveguide. Our scheme enriches the implementation of quantum routing and prompts the development of solid-state quantum networks.
@@ I Introduction
Quantum networks formed by interconnecting multiple quantum channels between quantum nodes serve as a significant constituent for the realization of quantum communication, quantum computing, and quantum sensing [1, 2, 3, 4, 5, 6, 7]. A central challenge in the development of practical quantum networks lies in achieving efficient internode quantum information transfer, a task in which quantum routers play an essential role [8, 9, 10, 11, 12]. Extensive research on quantum routing has been conducted on various physical platforms, such as superconducting circuits [13, 14, 15, 16, 17], cavity optomechanical systems [18, 19], atomic systems [20, 21, 22], the Su-Schrieffer-Heeger (SSH) model [23, 24], and cavity quantum electrodynamics (QED) [25]. In recent years, there has been notable growth in exploring quantum routing within waveguide QED systems [26, 27, 28, 29, 30, 31, 32].
@@
Diamonds have recently attracted considerable attention as a promising platform for solid-state quantum information processing due to their unique physical properties [33, 34, 35, 36, 37, 38, 39]. Besides the nitrogen-vacancy center in diamond [40, 41, 42, 43, 44, 45, 46, 47, 48], Si-vacancy (SiV) center in diamond has emerged as a leading platform among solid-state quantum systems [49, 50, 51]. The strain-induced interaction between SiV centers and the phonon modes achieves spin-phonon coupling strengths up to the megahertz scale, significantly exceeding the spin-photon coupling strength in nitrogen-vacancy center [52, 53, 54]. This characteristic offers us a novel approach to phonon-mediated quantum manipulation. Based on this advantage, numerous studies have been conducted, including the development of phonon-based quantum networks [55], the generation of spin-squeezed states [56, 57, 58, 59, 60], and quantum simulation of the SSH models [61, 62, 63]. The spin-phonon coupling properties and scalability of diamond waveguides endow the coupled system of SiV centers and phonons as a natural platform for solid-state quantum routers. However, two fundamental challenges remain in the realization of the quantum routing function [8]. First, remote entanglement is required as an essential resource in quantum routing. Second, the inevitable environmental noise in diamond waveguides causes decoherence on both quantum entanglement and state transfer of the SiV centers. Thus, robust preparation of long-range entanglement and effective suppression of decoherence during quantum-state transmission are critical for advancing quantum routing.
@@
In this work, in contrast to the conventional quantum-router schemes relying on active path-switching architectures [21, 64, 65, 66, 67, 68, 69], we propose a quantum router operating on a principle similar to that of a classical WiFi. Enabling a simultaneous transfer of quantum states from a single source node to multiple target nodes, our scheme works in the non-Markovian dynamics of an array of SiV centers interacting with the phonon modes in a one-dimensional diamond waveguide. We find that the strain-induced strong spin-phonon coupling in the diamond waveguide drives the hybrid SiV-phonon system to form multiple bound states, whose formation makes the long-distance entanglement among multiple nodes persistent and the multiplexed quantum-state transfer via the waveguide robust to the decoherence. Our result efficiently solves two key challenges in quantum routing. It not only enriches the implementation approaches of quantum routing, but also provides an effective pathway for scalable noise-resistant quantum networks.
@@ II System
where $V=Lab$ is the volume of the waveguide, $\varrho$ is the material density, and ${\bf u}_{n,k}^{\perp}(y,z)$ is the transverse profile of the displacement field, with $k$ being the wave vector and $n$ being the branch index. Then the Hamiltonian of the phonon is $\hat{H}_{\text{ph}}=\hbar\sum_{n,k}\omega_{n,k}\hat{a}_{n,k}^{\dagger}\hat{a}_{n,k}$, where $\hat{a}_{n,k}$ is the annihilation operator of the $k$th phonon mode with frequency $\omega_{n,k}$ [71, 72, 52]. The frequencies of the photons in the waveguides for a small cross section and a longe distance $L$ become continuous.
@@
where $|\hat{\sigma}_{j}=\ket{1}_{j}\bra{3}$ is the transition operator from $|3\rangle$ to $|1\rangle$ with the transition frequency $\omega_{0}$. The transition frequency is widely tunable from $46\,\mathrm{GHz}$ to $1.2\,\mathrm{THz}$ via static strain control [54]. The dispersion relation of the photons is taken to be linear as $\omega_{k}=vk$, where $v$ is the group velocity of the longitudinal phonon modes [60, 55].
@@ III Non-Markovian Quantum Routing
We prepare only the first SiV center in the excited state and all the phonon modes in the vacuum state initially, i.e., $|\Psi(0)\rangle=\hat{\sigma}_{1}^{\dagger}|G,\{0_{k}\}\rangle$. Here, $|G\rangle$ denotes that all the SiV centers are in their ground state and $|\{0_{k}\}\rangle$ is the vacuum state of the phonon modes in the waveguide. It is straightforward to verify that the total excitation number $\hat{\mathcal{N}}=\sum_{j}\hat{\sigma}_{j}^{\dagger}\hat{\sigma}_{j}+\sum_{k}\hat{a}_{k}^{\dagger}\hat{a}_{k}$ is conserved due to $\hat{H},\hat{\mathcal{N}}]=0$. Therefore, the time evolved state is expressed as
@@
where $h_{jl}(t-\tau)=\int_{0}^{\infty}d\omega J_{jl}(\omega)e^{-i\omega(t-\tau)}$ is the correlation function of the phonon mode and $J_{jl}(\omega)=2\pi\sum_{k}g_{jk}g_{lk}^{*}\delta(\omega-\omega_{k})$ denotes the correlated spectral density between the $j\text{th}$ and $l\text{th}$ SiV centers. The existence of such a correlation shows that the phonons in the diamond waveguide effectively mediates indirect couplings among the spatially separated SiV centers. This phonon-mediated coupling naturally implements an intrinsic quantum routing mechanism by establishing multiplexed coherent channels among distant SiV centers. The spectral density function is explicitly derived as
@@
where $\mathbf{c}(t)=\left(\begin{array}[]{ccccc}c_{1}(t)&\cdots&c_{N}(t)\\ \end{array}\right)^{\text{T}}$, $\mathbf{h}(t-\tau)=\int_{0}^{\infty}d\omega e^{-i\omega(t-\tau)}\mathbf{J}(\omega)$ is an NN-by-NN matrix with elements $h_{jl}(t-\tau)$, and $\mathbf{J}(\omega)$ is a spectral density matrix with elements $J_{jl}(\omega)$.
@@
where ${\bf Z}(t)=\sum_{j=1}^{N}\text{Res}[\tilde{\mathbf{c}}(-i\varpi_{j}^{b})]e^{-i\varpi_{j}^{b}t}$ and $\text{Res}[\tilde{\mathbf{c}}(-i\varpi_{j}^{b})]$ denotes the residue contributed by the $j\text{th}$ bound state. The second term originates from the continuous energy band. This term oscillates with a continuously varying frequency $\varpi$ and decays to zero in the long-time limit due to out-of-phase interference. Therefore, if no bound states exist, $\lim_{t\to\infty}\mathbf{c}(t)=\mathbf{0}$, indicating the complete decoherence and failure of quantum-information transfer. In contrast, the formation of bound states leads to $\lim_{t\to\infty}\mathbf{c}(t)=\mathbf{Z}(t)$, signifying the suppression of decoherence and a persistent quantum-information transfer among different SiV centers. Being absent in the Markov approximation, this behavior is a distinctive feature of non-Markovian dynamics. Being enabled by this mechanism, we realize a quantum router based on diamond waveguides. Analogous to conventional WiFi networks but operating through non-Markovian dynamics, this quantum router utilizes bound-state formation to establish stable entanglement among the SiV center nodes. As we will see later, it enables a parallel one-to-many quantum state transfer within the network.
@@
The bipartite entanglement between the SiV centers in a state $\rho$ is quantified by the concurrence $C=\max(0,\lambda_{1}-\lambda_{2}-\lambda_{3}-\lambda_{4})$, where $\lambda_{i}$ are the eigenvalues of the matrix $\rho(\hat{\sigma}_{y}\otimes\hat{\sigma}_{y})\rho^{*}(\hat{\sigma}_{y}\otimes\hat{\sigma}_{y})$ in the decreasing order [79]. In our case, we have
@@
Figure 2: (a) Energy spectrum of the total system obtained by numerically solving Eq. (9). (b) Time evolution of the concurrence $C(t)$ obtained by numerically solving Eq. (8) for different values of $\omega_{0}$. The orange, blue, and purple curves correspond to the cases with two, one, and zero bound states, respectively. The red dotted lines represent the analytical results of Eq. (14), showing a good agreement with the numerical results. We use $\omega_{c}=7\Delta$, $S=100\,\mathrm{nm}^{2}$, $d/2\pi\approx 4\,\mathrm{PHz}$, $\delta x\equiv|x_{1}-x_{2}|=10\,\mathrm{nm}$, $v\approx 1\times 10^{4}\,\text{m/s}$, and $N=2$.
@@
We begin by examining the entanglement characteristics of the quantum router under the non-Markovian dynamics. As depicted in Fig. 2, we show both the energy spectrum of the SiV-phonon system and the corresponding dynamics of concurrence $C(t)$ at different SiV transition frequencies $\omega_{0}$. We find that, in contrast to the complete decay to zero when $\omega_{0}=1.8\Delta$, $C(t)$ asymptotically approaches a finite value when $\omega_{0}=1.525\Delta$. It is surprising to see that $C(t)$ even tends to a lossless Rabi-like oscillation when $\omega_{0}=1.05\Delta$. Being completely absent in the conventional Markov approximate dynamics, these diverse behaviors are well explained by the features of the energy spectrum. We observe that the two branches of bound states in the band gap divide the spectrum into three distinct regions. When $\omega_{0}>1.58\Delta$, no bound state is formed and $C(t)$ decays entirely to zero. When $1.07\Delta<\omega_{0}\leq 1.58\Delta$, one bound state emerges, which leads to a finite steady-state entanglement. When $\omega_{0}\leq 1.07\Delta$, the formation of two bound states results in persistent Rabi-like oscillations of $C(t)$ in a frequency $|\varpi_{1}^{b}-\varpi_{2}^{b}|$. These diverse dynamical features excellently match the analytical predictions of Eq. (14). It demonstrates that the bound states and non-Markovian effects are crucial in sustaining entanglement for the quantum router. The result also reveals that, by appropriately selecting the operating frequency $\omega_{0}$ of SiV centers, persistent quantum correlation is achievable for the distant SiV centers even there is no direct interaction between them.
@@
To further test the robustness of entanglement and state transfer in the quantum router against increasing inter-SiV distance, Fig. 4 shows the energy spectrum, the steady-state entanglement $C(\infty)$, and the fidelity $F(\infty)$ as a function of the distance $\delta x$. When $\delta x<9\mathrm{nm}$, only one bound state exists and thus $C(\infty)$ stabilizes at a finite value and $F(\infty)$ exhibits a periodic oscillation with its maxima and minima matching with the values evaluated from Eq. (14). When $9\mathrm{nm}\leq\delta x<20\mathrm{nm}$, two bound states appear. It causes both of $C(\infty)$ and $F(\infty)$ to exhibit persistent oscillations. The one-to-one correspondence between the energy-spectrum features in Fig. 4(a) and the dynamical behaviors in Figs. 4(b) and 4(c) demonstrates that the bound states play a key role in establishing a persistent interconnent between the SiV centers even when their distance is large.
@@
Figure 5(a) shows three bound-state branches emerging in the energy spectrum of the three-node quantum router. In the absence of bound states, decoherence causes the entanglement between any bipartite partition to decay to zero and the fidelity between SiV1 and SiV2/SiV3 nodes to decay to its lower limit of $0.25$, see Figs. 5(b) and 5(c). When three bound states are present, SiV1 preserves simultaneous entanglement with SiV2 and SiV3 in the quantum router. The entanglement in this case exhibits tri-frequency oscillations, a behavior clearly demonstrated by the red solid and green dashed curves in Figs. 5(b) and 5(c). This configuration also significantly improves the state transfer fidelity from SiV1 to SiV2/SiV3, with its dynamics displaying complex multi-frequency oscillations. The exact match between analytical (red solid) and numerical (blue dashed) results in Figs. 5(b) and 5(c) conclusively validates the essential role of bound states in the quantum router. Similarly to the $N=2$ case, the three-SiV quantum router maintains quantum coherence over long distances and exhibits the spatial robustness. As shown in Fig. 6, the energy spectrum is presented along with the steady-state values of $C_{12}(\infty)/C_{13}(\infty)$ and $F_{12}(\infty)/F_{13}(\infty)$ as a function of the distance $\delta x$. The results demonstrate that, under the protection of bound states, both the entanglement and the fidelity in the quantum router can maintain either stable finite values or persistent multi-frequency oscillations even at long-distance node separations, revealing remarkable decoherence-resistant properties. These result demonstrates that the formation of bound states not only can sustain persistent entanglement among multiple nodes but also realize parallel one-to-many quantum state transfer. Thus, it establishes a robust noise-resistant quantum router for multi-node networks.

### v2

@@ Abstract
As a key component of quantum networks, the quantum router distributes quantum information among different quantum nodes. The silicon-vacancy (SiV) center in diamond offers a promising platform for quantum technology due to its strong strain-induced coupling with phonons. However, the development of a practical quantum router faces the challenges of achieving long-range entanglement and suppressing decoherence. Here, we propose a non-Markovian quantum router based on a diamond waveguide embedded with an array of SiV centers as the quantum nodes. Unlike conventional channel-switching methods, our design enables parallel quantum-state transfer from a single input node to multiple target nodes, analogous to a classical WiFi router. We demonstrate that persistent entanglement and suppressed decoherence of the SiV centers over long distances are achievable when bound states are present in the energy spectrum of the total system formed by the SiV centers and the phonon waveguide. Our scheme enriches the implementation of quantum routing and prompts the development of solid-state quantum networks.
@@ I Introduction
Quantum networks formed by interconnecting multiple quantum channels between quantum nodes serve as a significant constituent for the realization of quantum communication, quantum computing, and quantum sensing [1, 2, 3, 4, 5, 6, 7]. A central challenge in the development of practical quantum networks lies in achieving efficient internode quantum information transfer, a task in which quantum routers play an essential role [8, 9, 10, 11, 12]. Extensive research on quantum routing has been conducted on various physical platforms, such as superconducting circuits [13, 14, 15], cavity quantum electrodynamics (QED) systems [16, 17, 18], and atomic and optical systems [19, 20, 21]. While the above platforms have advanced quantum routing, they still face practical limitations such as on‑chip dissipation, miniaturization hurdles, and dependence on complex external controls. Recently, the outstanding performance of diamond-based color centers in building phonon networks has motivated us to explore the implementation of quantum routing on this promising platform [22, 23].
@@
Diamond color centers, such as the germanium-vacancy (GeV), nitrogen-vacancy (NV), tin-vacancy (SnV), and silicon-vacancy (SiV) systems, play significant roles in quantum science and technology [24, 25, 26, 27, 28, 29, 23, 30, 31]. Although NV centers are widely employed and exhibit outstanding coherence properties [32, 33, 34, 35, 36, 37, 38, 39, 40], their intrinsically weak coupling to diamond phonon modes limits their performance in strong-coupling regimes. In contrast, centers such as SiV and SnV feature strain-mediated phonon couplings that exceed those of NV centers by several orders of magnitude [41, 42, 43, 23], establishing them as leading candidates among solid-state quantum platforms [44, 45, 46, 23]. This characteristic offers us a different approach to phonon-mediated quantum manipulation. Based on this advantage, numerous studies have been conducted, including the development of phonon-based quantum networks [22], the generation of spin-squeezed states [47, 48, 49, 50, 51], and quantum simulation of the Su-Schrieffer-Heeger models [52, 53, 54]. The spin-phonon coupling properties and scalability of diamond waveguides endow the coupled system of SiV centers and phonons as a natural platform for solid-state quantum routers. However, three fundamental challenges remain in the realization of the quantum routing function [8]. First, the generated entanglement decreases dramatically with increasing the distance among the quantum nodes. Second, the inevitable environmental noise in diamond waveguides causes decoherence on both quantum entanglement and state transfer of the SiV centers. Third, conventional quantum routers typically operate as end-to-end switches [55, 56, 57, 58, 59, 60, 61], which makes it hard to support a parallel state transfer to multiple nodes. Thus, achieving robust long-range entanglement, suppressing decoherence during state transmission, and overcoming the end‑to‑end limitation of conventional quantum routers are highly desired.
@@
In this work, to overcome these limitations, we propose a quantum router operating on a principle similar to that of a classical WiFi. Enabling a simultaneous transfer of quantum states from a single source node to multiple target nodes, our scheme works in the non-Markovian dynamics of an array of SiV centers interacting with the phonon modes in a one-dimensional diamond waveguide. We find that the strain-induced strong spin-phonon coupling in the diamond waveguide drives the hybrid SiV-phonon system to form multiple bound states, whose formation makes the long-distance entanglement among multiple nodes persistent and the multiplexed quantum-state transfer via the waveguide robust to the decoherence. Our result efficiently solves these key challenges in quantum routing. It not only enriches the implementation approaches of quantum routing, but also provides an effective pathway for scalable noise-resistant quantum networks.
@@ II System
where $V=Lab$ is the volume of the waveguide, $\varrho$ is the material density, and ${\bf u}_{n,k}^{\perp}(y,z)$ is the transverse profile of the displacement field, with $k$ being the wave vector and $n$ being the branch index. Then the Hamiltonian of the phonon is $\hat{H}_{\text{ph}}=\hbar\sum_{n,k}\omega_{n,k}\hat{a}_{n,k}^{\dagger}\hat{a}_{n,k}$, where $\hat{a}_{n,k}$ is the annihilation operator of the $k$th phonon mode with frequency $\omega_{n,k}$ [63, 64, 41]. The frequencies of the phonons in the waveguides for a small cross section and a longe distance $L$ become continuous.
@@
where $|\hat{\sigma}_{j}=\ket{1}_{j}\bra{3}$ is the transition operator from $|3\rangle$ to $|1\rangle$ with the transition frequency $\omega_{0}$. The transition frequency is widely tunable from $46\,\mathrm{GHz}$ to $1.2\,\mathrm{THz}$ via static strain control [43]. The dispersion relation of the phonons is taken to be linear as $\omega_{k}=vk$, where $v$ is the group velocity of the longitudinal phonon modes [51, 22].
@@ III Non-Markovian Quantum Routing
We prepare only the first SiV center in the excited state and all the phonon modes in the vacuum state initially, i.e., $|\Psi(0)\rangle=\hat{\sigma}_{1}^{\dagger}|G,\{0_{k}\}\rangle$. Here, $|G\rangle$ denotes that all the SiV centers are in their ground state and $|\{0_{k}\}\rangle$ is the vacuum state of the phonon modes in the waveguide. This vacuum-state assumption is physically well justified in the low-temperature regime $T\ll\hbar\Delta/k_{B}\approx 2.2\mathrm{~K}$, where the thermal occupation number is entirely negligible [22]. It is straightforward to verify that the total excitation number $\hat{\mathcal{N}}=\sum_{j}\hat{\sigma}_{j}^{\dagger}\hat{\sigma}_{j}+\sum_{k}\hat{a}_{k}^{\dagger}\hat{a}_{k}$ is conserved due to $\hat{H},\hat{\mathcal{N}}]=0$. Therefore, the time evolved state is expressed as
@@
where $h_{jl}(t-\tau)=\int_{0}^{\infty}d\omega J_{jl}(\omega)e^{-i\omega(t-\tau)}$ is the correlation function of the phonon mode and $J_{jl}(\omega)=2\pi\sum_{k}g_{jk}g_{lk}^{*}\delta(\omega-\omega_{k})$ denotes the correlated spectral density between the $j\text{th}$ and $l\text{th}$ SiV centers. The existence of such a correlation shows that the phonons in the diamond waveguide effectively mediate indirect couplings among the spatially separated SiV centers. This phonon-mediated coupling naturally implements an intrinsic quantum routing mechanism by establishing multiplexed coherent channels among distant SiV centers. The spectral density function is explicitly derived as
@@
where $\mathbf{c}(t)=\left(\begin{array}[]{ccccc}c_{1}(t)&\cdots&c_{N}(t)\\ \end{array}\right)^{\text{T}}$, $\mathbf{h}(t-\tau)=\int_{0}^{\infty}d\omega e^{-i\omega(t-\tau)}\mathbf{J}(\omega)$ is an $N\times N$ matrix with elements $h_{jl}(t-\tau)$, and $\mathbf{J}(\omega)$ is a spectral density matrix with elements $J_{jl}(\omega)$.
@@
where ${\bf Z}(t)=\sum_{j=1}^{N}\text{Res}[\tilde{\mathbf{c}}(-i\varpi_{j}^{b})]e^{-i\varpi_{j}^{b}t}$ and $\text{Res}[\tilde{\mathbf{c}}(-i\varpi_{j}^{b})]$ denotes the residue contributed by the $j\text{th}$ bound state. The second term originates from the continuous-energy band. This term oscillates with a continuously varying frequency $\varpi$ and decays to zero in the long-time limit due to out-of-phase interference. Therefore, if no bound states exist, $\lim_{t\to\infty}\mathbf{c}(t)=\mathbf{0}$, indicating the complete decoherence and failure of quantum-information transfer. In contrast, the formation of bound states leads to $\lim_{t\to\infty}\mathbf{c}(t)=\mathbf{Z}(t)$, signifying the suppression of decoherence and a persistent quantum-information transfer among different SiV centers. Being absent in the Markov approximation, this behavior is a distinctive feature of non-Markovian dynamics. Being enabled by this mechanism, we realize a quantum router based on diamond waveguides. Analogous to conventional WiFi networks but operating through non-Markovian dynamics, this quantum router utilizes bound-state formation to establish stable entanglement among the SiV center nodes. As we will see later, it enables a parallel one-to-many quantum state transfer within the network.
@@
Figure 2: (a) Energy spectrum of the total system obtained by numerically solving Eq. (9). (b) Time evolution of the concurrence $C(t)$ obtained by numerically solving Eq. (8) for different values of $\omega_{0}$. The orange, blue, and purple curves correspond to the cases with two, one, and zero bound states, respectively. The red dotted lines represent the analytical results of Eq. (14), showing a good agreement with the numerical results. (c) Evolution of concurrence $C(t)$ when a random fluctuation $\chi$ is added on the frequency $\omega_{0}=\Delta$ of the SiV centers. The red and green dashed lines are the average values obtained by averaging over 100 random configurations when $\chi=[-0.5\Delta,0.5\Delta]$ and $\Delta,5\Delta]$. The purple and orange regions are their respective variations. The blue solid lines are the fluctuation-free results. We use $\omega_{c}=7\Delta$, $S=100\,\mathrm{nm}^{2}$, $d/2\pi\approx 4\,\mathrm{PHz}$, $\delta x\equiv|x_{1}-x_{2}|=10\,\mathrm{nm}$, $v\approx 1\times 10^{4}\,\text{m/s}$, and $N=2$.
The bipartite entanglement between the SiV centers in a state $\rho$ is quantified by the concurrence $C=\max(0,\lambda_{1}-\lambda_{2}-\lambda_{3}-\lambda_{4})$, where $\lambda_{i}$ are the square roots of the eigenvalues of the matrix $\rho(\hat{\sigma}_{y}\otimes\hat{\sigma}_{y})\rho^{*}(\hat{\sigma}_{y}\otimes\hat{\sigma}_{y})$ in the decreasing order [71]. In our case, we have
@@
We begin by examining the entanglement characteristics of the quantum router under the non-Markovian dynamics. As depicted in Fig. 2, we show both the energy spectrum of the SiV-phonon system and the corresponding dynamics of concurrence $C(t)$ at different SiV transition frequencies $\omega_{0}$. We find that in contrast to the complete decay to zero when $\omega_{0}=1.8\Delta$, $C(t)$ asymptotically approaches a finite value when $\omega_{0}=1.525\Delta$. It is surprising to see that $C(t)$ even tends to a lossless Rabi-like oscillation when $\omega_{0}=1.05\Delta$. Being completely absent in the conventional Markov approximate dynamics, these diverse behaviors are well explained by the features of the energy spectrum. We observe that the two branches of bound states in the band gap divide the spectrum into three distinct regions. When $\omega_{0}>1.58\Delta$, no bound state is formed and $C(t)$ decays entirely to zero. When $1.07\Delta<\omega_{0}\leq 1.58\Delta$, one bound state emerges, which leads to a finite steady-state entanglement. When $\omega_{0}\leq 1.07\Delta$, the formation of two bound states results in persistent Rabi-like oscillations of $C(t)$ in a frequency proportional to $|\varpi_{1}^{b}-\varpi_{2}^{b}|$. These diverse dynamical features excellently match the analytical predictions of Eq. (14). It demonstrates that the bound states and non-Markovian effects are crucial in sustaining entanglement for the quantum router. The result also reveals that by appropriately selecting the operating frequency $\omega_{0}$ of SiV centers, persistent quantum correlation is achievable for the distant SiV centers even though there is no direct interaction between them.
Our analysis has assumed identical SiV centers for the sake of clarity. However, practical devices inevitably exhibit a random distribution in their quantum properties. To assess the impact of the random distribution of the frequency of the $N$ SiV centers on our quantum-router scheme, we have performed a numerical calculation by adding a random frequency shift $\chi$ on the frequency of the SiV centers. As shown in Fig. (2)(c), when the random frequency fluctuation is confined to a narrow range $0.5\Delta,0.5\Delta]$, the averaged concurrence $C(t)$ remains close to the ideal case, indicating that the router can still function well. However, when the range of the frequency fluctuation increases to $\Delta,5\Delta]$, the averaged $C(t)$ drops significantly and the fluctuations become large, implying a severe degradation in both entanglement generation and routing fidelity. Furthermore, when the $N$ SiV centers possess different frequencies, the quantitative condition for forming the bound states might be changed. However, as long as the bound states are present, the persistent entanglement between the SiV centers over long distances is achievable. These indicate that there is a clear tolerance to the parameter randomness for the reliable operation of our quantum-router scheme. The extremely large frequency expansion and mismatch of SiV centers would definitely lead to a decline in entanglement-generation efficiency, a reduction in quantum-state fidelity, and an increase in the unpredictability of system behavior. From an experimental perspective, it may be feasible to suppress these unwanted imperfections by exploiting static strain control to tune the frequencies of SiV centers [43].
@@
To further test the robustness of entanglement and state transfer in the quantum router against increasing inter-SiV distance, Fig. 4 shows the energy spectrum, the steady-state entanglement $C(\infty)$, and the fidelity $F(\infty)$ as a function of the distance $\delta x$. When $\delta x<9\mathrm{nm}$, only one bound state exists and thus $C(\infty)$ stabilizes at a finite value and $F(\infty)$ exhibits a periodic oscillation with its maxima and minima matching with the values evaluated from Eq. (14). When $9\mathrm{nm}\leq\delta x<20\mathrm{nm}$, two bound states appear. It causes both $C(\infty)$ and $F(\infty)$ to exhibit persistent oscillations. The one-to-one correspondence between the energy-spectrum features in Fig. 4(a) and the dynamical behaviors in Figs. 4(b) and 4(c) demonstrates that the bound states play a key role in establishing a persistent interconnect between the SiV centers even when their distance is large. With further increasing the distance between the SiV centers, the eigenenergies $\hbar\varpi_{i}^{b}$ of the two bound states increasingly nearer. It causes the decrease of the oscillation frequency of the concurrence, which is proportional to $|\varpi_{1}^{b}-\varpi_{2}^{b}|$. However, our result that a persistent entanglement and suppressed decoherence of the SiV centers are achievable when two bound states are present in the energy spectrum of the total system formed by the SiV centers and the phonon waveguide does not change. This demonstrates the robustness of our quantum-router scheme to the separation of the SiV centers.
@@
Figure 5(a) shows three bound-state branches emerging in the energy spectrum of the three-node quantum router. In the absence of bound states, decoherence causes the entanglement between any bipartite partition to decay to zero and the fidelity between SiV1 and SiV2/SiV3 nodes to decay to its lower limit of $0.25$; see Figs. 5(b) and 5(c). When three bound states are present, SiV1 preserves simultaneous entanglement with SiV2 and SiV3 in the quantum router. The entanglement in this case exhibits trifrequency oscillations, a behavior clearly demonstrated by the red solid and green dashed curves in Figs. 5(b) and 5(c). This configuration also significantly improves the state-transfer fidelity from SiV1 to SiV2/SiV3, with its dynamics displaying complex multifrequency oscillations. The exact match between the analytical (red solid) and numerical (blue dashed) results in Figs. 5(b) and 5(c) conclusively validates the essential role of bound states in the quantum router. Similarly to the $N=2$ case, the three-SiV quantum router maintains quantum coherence over long distances and exhibits the spatial robustness. As shown in Fig. 6, the energy spectrum is presented along with the steady-state values of $C_{12}(\infty)/C_{13}(\infty)$ and $F_{12}(\infty)/F_{13}(\infty)$ as a function of the distance $\delta x$. The results demonstrate that under the protection of bound states, both the entanglement and the fidelity in the quantum router can maintain either stable finite values or persistent multifrequency oscillations even at long-distance node separations, revealing remarkable decoherence-resistant properties. These results demonstrate that the formation of bound states not only can sustain persistent entanglement among multiple nodes but also realize parallel one-to-many quantum-state transfer. Thus, it establishes a robust noise-resistant quantum router for multinode networks.